OpenStudy (anonymous):

The point (3,6) is on the graph of y=2 f(x+1)-4. Find the original point on the graph of y=f(x). I'm honestly at a lost and have no idea what to do.

OpenStudy (aravindg):

(3,6) satisfies y=2 f(x+1)-4 .So plug it in y=2 f(x+1)-4 .What do you get?

OpenStudy (anonymous):

so [y=2(3+1)-4 ]

OpenStudy (anonymous):

and y=2(6+1)-4?

OpenStudy (aravindg):

Nope

OpenStudy (anonymous):

The answer in the book is 4,5

OpenStudy (aravindg):

Put x=3 and y=6 at the same time

OpenStudy (anonymous):

then i get 6=4

OpenStudy (aravindg):

y=2 f(x+1)-4 6=2(f(3+1)-4 10=2f(4) 5=f(4) How you get 6=4?

OpenStudy (anonymous):

Ah, i see, I forgot that the numbers in the brackets manipulated x and the numbers outside the brackets manipulated y. Thank you!

OpenStudy (aravindg):

yw :)

OpenStudy (aravindg):

And Welcome to Openstudy!

OpenStudy (anonymous):

thanks